# Stochastic Control Methods for Optimization

### v2

@@ Abstract
In this work, we investigate a stochastic control framework for global optimization over both Euclidean spaces and the Wasserstein space of probability measures, where the objective function may be non-convex and/or non-differentiable. In the Euclidean setting, the original minimization problem is approximated by a family of regularized stochastic control problems; using dynamic programming, we analyze the associated Hamilton–Jacobi–Bellman equations and obtain tractable representations via the Cole–Hopf transformation and the Feynman–Kac formula. For optimization over probability measures, we formulate a regularized mean-field control problem characterized by a master equation, and further approximate it by controlled $N$-particle systems. We establish that, as the regularization parameter tends to zero (and as the particle number tends to infinity for the optimization over probability measures), the value of the control problem converges to the global minimum of the original objective. Building on the resulting probabilistic representations, Monte Carlo-based numerical schemes are proposed and numerical experiments are reported to illustrate the effectiveness of the methods and to support the theoretical convergence rates.
@@
Keywords: Optimization, mean-field control, Brownian bridge, Schrödinger bridge, Cole-Hopf transformation, mean-field game
@@ 1 Introduction
where $\mathcal{X}$ is a finite-dimensional space $\mathbb{R}^{d}$ or the space of probability measures and the objective function $G:\mathcal{X}\rightarrow\mathbb{R}$ may be non-convex and/or non-differentiable. Optimization problems of the form (1.1) arise in a wide range of applications, including operational research, machine learning, artificial intelligence, finance, engineering design, economics, and operations research. Traditional optimization methods, such as gradient descent and Newton’s method, often struggle with non-convex or non-differentiable functions due to the presence of multiple local minimas and the lack of gradient information; see [22, 59]. To address these challenges, we propose and study stochastic control methods to solve the optimization problem (1.1).
@@ 1.2.1 The master equation and $N$-Particle Approximation
For any random variable $\zeta$, denote by $\mathcal{L}(\zeta)$ its law/distribution. We define the cost functional
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
Substituting the optimal control (1.25) into the controlled SDEs (1.22), we obtain the SDEs for the optimal state processes $\{X_{t}^{i,0,x_{0};\theta_{i}^{*}}\}_{i=1}^{N}$ which at time 1 gives the approximation of the optimization problem (1.16). We establish the convergence of the proposed stochastic control method to the global minimizer of the functional $G$. Our main theoretical result establishes that the value computed via the $N$-particle approximations converges to the true global minimum of $G$ as $N\rightarrow\infty$ and $\varepsilon\rightarrow 0$.
@@
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C_{2}\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right)}_{\text{Regularization Error}},$ (1.26)
where $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and $C_{2}$ depends on $d$, the Lipschitz constant of $G$, and the Wasserstein-2 distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between $\delta_{x_{0}}$ and $\mu^{*}$.
For clarity of exposition, we assume that all particles start from the same deterministic point $x_{0}\in\mathbb{R}^{d}$. More generally, if the initial positions are sampled i.i.d. according to an arbitrary law $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then the total approximation error contains an additional term that accounts for the discrepancy between $\mu_{0}$ and the empirical initial distribution; see Corollary 3.4. Moreover, under the regularity and convexity conditions of Proposition 3.2, the constant $L_{\varepsilon}$ in the particle approximation estimate can be chosen uniformly in $\varepsilon$; see Theorem 3.3.
@@ 1.3 Literature, contributions, and organization of the paper
  • Numerical algorithms are proposed based on the probabilistic representations of the value functions and optimal controls. Numerical experiments are conducted to validate the theoretical findings and demonstrate the effectiveness of the proposed methods. To reproduce the results, the code is available at
@@
The remainder of this paper is structured as follows. Section 2 develops the stochastic control formulation for global optimization in Euclidean spaces, provides a detailed convergence analysis as the regularization parameter $\varepsilon\to 0$, establishes the well-posedness of the associated control problems, and reports numerical experiments. Section 3 extends the proposed framework to optimization over the Wasserstein space of probability measures, presenting the mean-field control formulation, deriving the corresponding convergence results (including the $N$-particle approximation), and illustrating the methodology through numerical studies. Finally, the proofs for Proposition 3.2 and Theorem 3.3 are provided in Appendix A and Appendix B, respectively.
@@ 2.3 Numerical approximations with examples
It is worth noting that as our primary emphasis is on the theoretical development and analysis of the proposed stochastic control method, the numerical experiments above are presented solely for illustrative purposes, and the reported parameter choices are not optimized for accuracy or computational efficiency. In practical implementations, in addition to key parameters such as the regularization level $\varepsilon$, the particle number $N$, the Monte Carlo sample size $S$, the number of outer iterations $L$, the coupling weight $\lambda$, and the time discretization level $M$, one may further tune the time horizon $T$, employ more sophisticated Monte Carlo estimators or variance-reduction techniques, and introduce suitable scalings of the driving Wiener process $W$. A comprehensive numerical investigation, including systematic parameter calibration and detailed comparisons with established global optimization algorithms, is left for future work.
@@ 3 Optimization over probability measures: $\mathcal{X}=\mathcal{P}_{2}(\mathbb{R}^{d})$
This section is devoted to the stochastic control method introduced in Section 1.2 for global optimization over the space of probability measures. We provide a convergence analysis of the proposed particle approximation scheme, including quantitative bounds for both the finite-particle approximation error and the regularization error.
@@ 3.1 Finite Particle Approximation
To extend the result to a general probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, one may rely on the Lipschitz continuity of the value function $\mathcal{V}_{\varepsilon}$; note that we do not need such an extension in Theorem 1.2 as we choose all particles starting from a fixed point. In fact, the value function $\mathcal{V}_{\varepsilon}$ is Lipschitz continuous with respect to the Wasserstein-2 distance under the Lipschitz assumptions on $G$. For any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, by the definition of $\mathcal{V}_{\varepsilon}$, we have
@@ 3.2 Regularization error
Let $G:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be Hölder continuous with respect to the Wasserstein-2 distance, i.e., there exist $L_{G}>0$ and $\alpha\in(0,1]$ such that $|G(\mu)-G(\nu)|\leq L_{G}|\mathcal{W}_{2}(\mu,\nu)|^{\alpha}$ for all $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Assume that $G$ admits a global minimizer $\mu^{*}$ with a finite second moment (i.e., $\int|x|^{2}d\mu^{*}(x)<\infty$). Let $\mathcal{V}_{0}=\inf_{\mu}G(\mu)$. Then the value function $\mathcal{V}_{\varepsilon}$ satisfies
@@
for some constant $C>0$ depending on $d$, $L_{G}$, $\alpha$, and the Wasserstein-2 distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between the Dirac measure at $x_{0}$ and $\mu^{*}$.
@@
As $G$ is Hölder continuous with respect to $\mathcal{W}_{2}$, we have $|G(\mu^{\lambda})-G(\mu^{*})|\leq L_{G}\cdot|\mathcal{W}_{2}(\mu^{\lambda},\mu^{*})|^{\alpha}.$ Recalling that the measure $\mu^{\lambda}$ is obtained by adding independent noise $\sqrt{\lambda}Z$ to $\mu^{*}$, with $Z\sim\mathcal{N}(0,I_{d})$. The Wasserstein-2 distance is bounded by the second moment of the noise. i.e.,
@@ 3.3 Numerical approximation with examples
For this purely interaction-driven energy, the global minimizer is not unique but known explicitly: it is the uniform probability measure supported on a unit circle, with minimal energy equal to $0.75$, commonly referred to as the circle law [14]. Starting from the initial distribution concentrated at the origin, the proposed stochastic control method drives the particle system to expand radially and, by terminal time, produces an empirical distribution that closely approximates the uniform measure on the unit circle.
@@
In the above 2D Newtonian swarm example, it is not appropriate to study the sensitivity with respect to the regularization parameter $\varepsilon$ because the particle approximation error also depends on $\varepsilon$; see Theorem 3.3. To isolate the regularization error, we consider the following example that satisfies the regularity and convexity assumptions in Proposition 3.2 under which the constant $L_{\varepsilon}$ in Thoerem 3.3 is independent of $\varepsilon$.
@@
$G\mu)={\int_{\mathbb{R}^{2}}V(x)\,d\mu(x)}+{\frac{1}{2}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\Phi(x-y)\,d\mu(x)\,d\mu(y)},$ (3.17)
@@
for $c_{L}=(-2,0)$, $c_{R}=(2,0)$, and $R=1$. The confinement potential $V$ has two wells centered at $c_{L}$ and $c_{R}$, each shaped like a hula hoop of radius $R$. The interaction potential $\Phi$ is the Newtonian potential in 2D, which encourages particles to spread out. By theories of equilibrium measures in [55], the minimizer forms two disjoint, thickened annular bands with density increasing from the inner to the outer rim, achieving a global minimum energy where the outward pressure of logarithmic repulsion exactly balances the inward force of the quartic confinement; as the confinement potential is not smooth, it is not straightforward to derive an explicit formula for the density profile. In the numerical experiment reported in Figure 6, we take $N=400$ particles and discretize the time interval $[0,1]$ with $M=1000$ steps. The regularization parameter is set to $\varepsilon=1e-299$, and the conditional expectations in (3.11) are approximated using $S=100$ Monte Carlo samples. We perform a single outer iteration ($L=1$) and initialize all particles with standard normal distribution $N(0,1)$ on both the x and y axes. The proposed stochastic control method successfully drives the particle system to approximate the minimizer supported around the two circles by terminal time.

### v3

# arXiv:2601.01248v3[math.OC] 02 Mar 2026

@@ Abstract
In this work, we investigate a stochastic control framework for global optimization over both Euclidean spaces and the Wasserstein space of probability measures, where the objective function may be non-convex and/or non-differentiable. In the Euclidean setting, the original minimization problem is approximated by a family of regularized stochastic control problems; using dynamic programming, we analyze the associated Hamilton–Jacobi–Bellman equations and obtain tractable representations via the Cole–Hopf transformation and the Feynman–Kac formula. For optimization over probability measures, we formulate a regularized mean-field control problem characterized by a master equation, and further approximate it by controlled $N$-particle systems. We establish that, as the regularization parameter tends to zero (and as the particle number tends to infinity for the optimization over probability measures), the value of the control problem converges to the global minimum of the original objective. Building on the resulting probabilistic representations, we propose the Monte Carlo-based numerical schemes that are derivative-free due to the utilization of the Bismut-Elworthy-Li formula and numerical experiments are reported to illustrate the effectiveness of the methods and to support the theoretical convergence rates.
@@
Keywords: Optimization, mean-field control, Brownian bridge, Schrödinger bridge, Cole-Hopf transformation, generative modeling
@@ 1 Introduction
where $\mathcal{X}$ is a finite-dimensional space $\mathbb{R}^{d}$ or the space of probability measures and the objective function $G:\mathcal{X}\rightarrow\mathbb{R}$ may be non-convex and/or non-differentiable. Optimization problems of the form (1.1) arise in a wide range of applications, including operational research, machine learning, artificial intelligence, finance, engineering design, economics, and operations research. Traditional optimization methods, such as gradient descent and Newton’s method, often struggle with non-convex or non-differentiable functions due to the presence of multiple local minimas and the lack of gradient information; see [22, 61]. To address these challenges, we propose and study stochastic control methods (SCM) to solve the optimization problem (1.1).
@@ 1.2.1 The master equation and $N$-Particle Approximation
For any square-integrable random variable $\zeta$, denote by $\mathcal{L}(\zeta)$ its law/distribution. We define the cost functional
@@ 1.2.2 Main result on convergence analysis as $\varepsilon\to 0$ and $N\to\infty$
Substituting the optimal control (1.25) into the controlled SDEs (1.22), we obtain the SDEs for the optimal state processes $\{X_{t}^{i,0,x_{0};\theta_{i}^{*}}\}_{i=1}^{N}$ which at time 1 gives the approximation of the optimization problem (1.16). We establish the convergence of the proposed SCM to the global minimizer of the functional $G$. Our main theoretical result establishes that the value computed via the $N$-particle approximations converges to the true global minimum of $G$ as $N\rightarrow\infty$ and $\varepsilon\rightarrow 0$.
@@
$\left|\frac{1}{N}v^{N}_{\varepsilon}(0,\mathbf{x}^{0})-\mathcal{V}_{0}\right|\leq\underbrace{\frac{L_{\varepsilon}}{2N}}_{\text{Particle Error}}+\underbrace{C_{2}\cdot\varepsilon\ln\left(\frac{1}{\varepsilon}\right)}_{\text{Regularization Error}},$ (1.26)
where $\mathbf{x}^{0}=(x^{0}_{1},\ldots,x^{0}_{N})=(x_{0},\ldots,x_{0})\in\mathbb{R}^{Nd}$, $L_{\varepsilon}$ depends on $\varepsilon$ and the regularity of $G$, and $C_{2}$ depends on $d$, the Lipschitz constant of $G$, and the 2-Wasserstein distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between $\delta_{x_{0}}$ and $\mu^{*}$.
The above theorem summarizes the convergence results presented in Theorems 3.3 and 3.5. For clarity of exposition, we assume that all particles start from the same deterministic point $x_{0}\in\mathbb{R}^{d}$. More generally, if the initial positions are sampled i.i.d. according to an arbitrary law $\mu_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, then the total approximation error contains an additional term that accounts for the discrepancy between $\mu_{0}$ and the empirical initial distribution; see Corollary 3.4. Moreover, under the regularity, convexity, and coercivity conditions of Proposition 3.2, the constant $L_{\varepsilon}$ in the particle approximation estimate can be chosen uniformly in $\varepsilon$; see Theorem 3.3.
@@ 1.3 Literature, contributions, and organization of the paper
  • Based on the probabilistic representations of the value functions and optimal controls, we propose the numerical algorithms that are derivative-free due to the utilization of the Bismut-Elworthy-Li formula. Numerical experiments are conducted to validate the theoretical findings and demonstrate the effectiveness of the proposed methods. To reproduce the results, the code is available at
@@
The remainder of this paper is structured as follows. Section 2 develops the stochastic control formulation for global optimization in Euclidean spaces, provides a detailed convergence analysis as the regularization parameter $\varepsilon\to 0$, establishes the well-posedness of the associated control problems, and reports numerical experiments. Section 3 extends the proposed framework to optimization over the Wasserstein space of probability measures, presenting the mean-field control formulation, deriving the corresponding convergence results (including the $N$-particle approximation), and illustrating the methodology through numerical studies. Then the conclusion in Section 4 summarizes the findings and discusses potential avenues for future research. Finally, the proofs for Proposition 3.2 and Theorem 3.3 are provided in Appendix A and Appendix B, respectively.
@@ 2.3 Numerical approximations with examples
It is worth noting that as our primary emphasis is on the theoretical development and analysis of the proposed SCM for global optimization, the numerical experiments above are presented solely for illustrative purposes, and the reported parameter choices are not optimized for accuracy or computational efficiency. In practical implementations, in addition to key parameters such as the regularization level $\varepsilon$, the particle number $N$, the Monte Carlo sample size $S$, the number of outer iterations $L$, the coupling weight $\lambda$, and the time discretization level $M$, one may further tune the time horizon $T$, employ more sophisticated Monte Carlo estimators or variance-reduction techniques, and introduce suitable scalings of the driving Wiener process $W$. A comprehensive numerical investigation, including systematic parameter calibration and detailed comparisons with established global optimization algorithms, is left for future work.
@@ 3 Optimization over probability measures: $\mathcal{X}=\mathcal{P}_{2}(\mathbb{R}^{d})$
This section is devoted to the SCM introduced in Section 1.2 for global optimization over the space of probability measures. We provide a convergence analysis of the proposed particle approximation scheme, including quantitative bounds for both the finite-particle approximation error and the regularization error.
@@ 3.1 Finite Particle Approximation
To extend the result to a general probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, one may rely on the Lipschitz continuity of the value function $\mathcal{V}_{\varepsilon}$; note that we do not need such an extension in Theorem 1.2 as we choose all particles starting from a fixed point. In fact, the value function $\mathcal{V}_{\varepsilon}$ is Lipschitz continuous with respect to the 2-Wasserstein distance under the Lipschitz assumptions on $G$. For any $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$, by the definition of $\mathcal{V}_{\varepsilon}$, we have
@@ 3.2 Regularization error
Let $G:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be Hölder continuous with respect to the 2-Wasserstein distance, i.e., there exist $L_{G}>0$ and $\alpha\in(0,1]$ such that $|G(\mu)-G(\nu)|\leq L_{G}|\mathcal{W}_{2}(\mu,\nu)|^{\alpha}$ for all $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$. Assume that $G$ admits a global minimizer $\mu^{*}$ with a finite second moment (i.e., $\int|x|^{2}d\mu^{*}(x)<\infty$). Let $\mathcal{V}_{0}=\inf_{\mu}G(\mu)$. Then the value function $\mathcal{V}_{\varepsilon}$ satisfies
@@
for some constant $C>0$ depending on $d$, $L_{G}$, $\alpha$, and the 2-Wasserstein distance ($\mathcal{W}_{2}(\delta_{x_{0}},\mu^{*})$) between the Dirac measure at $x_{0}$ and $\mu^{*}$.
@@
As $G$ is Hölder continuous with respect to $\mathcal{W}_{2}$, we have $|G(\mu^{\lambda})-G(\mu^{*})|\leq L_{G}\cdot|\mathcal{W}_{2}(\mu^{\lambda},\mu^{*})|^{\alpha}.$ Recalling that the measure $\mu^{\lambda}$ is obtained by adding independent noise $\sqrt{\lambda}Z$ to $\mu^{*}$, with $Z\sim\mathcal{N}(0,I_{d})$. The 2-Wasserstein distance is bounded by the second moment of the noise. i.e.,
@@ 3.3 Numerical approximation with examples
For this purely interaction-driven energy, the global minimizer is not unique but known explicitly: it is the uniform probability measure supported on a unit circle, with minimal energy equal to $0.75$, commonly referred to as the circle law [14]. Starting from the initial distribution concentrated at the origin, the proposed SCM drives the particle system to expand radially and, by terminal time, produces an empirical distribution that closely approximates the uniform measure on the unit circle.
@@
In the above 2D Newtonian swarm example, it is not appropriate to study the sensitivity with respect to the regularization parameter $\varepsilon$ because the particle approximation error also depends on $\varepsilon$; see Theorem 3.3. To isolate the regularization error, we consider the following example that satisfies the regularity, convexity, and coercivity assumptions in Proposition 3.2 under which the constant $L_{\varepsilon}$ in Thoerem 3.3 is independent of $\varepsilon$.
@@
$G(\mu)={\int_{\mathbb{R}^{2}}V(x)\,d\mu(x)}+{\frac{1}{2}\iint_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\Phi(x-y)\,d\mu(x)\,d\mu(y)},$ (3.17)
@@
for $c_{L}=(-2,0)$, $c_{R}=(2,0)$, and $R=1$. The confinement potential $V$ has two wells centered at $c_{L}$ and $c_{R}$, each shaped like a hula hoop of radius $R$. The interaction potential $\Phi$ is the Newtonian potential in 2D, which encourages particles to spread out. By theories of equilibrium measures in [57], the minimizer forms two disjoint, thickened annular bands with density increasing from the inner to the outer rim, achieving a global minimum energy where the outward pressure of logarithmic repulsion exactly balances the inward force of the quartic confinement; as the confinement potential is not smooth, it is not straightforward to derive an explicit formula for the density profile. In the numerical experiment reported in Figure 6, we take $N=400$ particles and discretize the time interval $[0,1]$ with $M=1000$ steps. The regularization parameter is set to $\varepsilon=10^{-299}$, and the conditional expectations in (3.11) are approximated using $S=100$ Monte Carlo samples. We perform a single outer iteration ($L=1$) and initialize all particles with standard normal distribution $N(0,1)$ on both the x and y axes. The proposed SCM successfully drives the particle system to approximate the minimizer supported around the two circles by terminal time.
@@
The above example further inpires us to apply the SCM to sampling and generative problems. In generative modeling and aritificial intelligence (AI), a fundamental problem is: given an empirical dataset $\mathbf{A}=\{x^{1},\dots,x^{J}\}\subset\mathbb{R}^{d}$ which defines an underlying distribution $\hat{\mu}=\frac{1}{J}\sum_{j=1}^{J}\delta_{x^{j}}$, how can one generate a new set of data $\mathbf{D}=\{y^{1},\dots,y^{N}\}$ such that the distribution of $\mathbf{D}$ is equal to or close to that of $\mathbf{A}$? This is nothing but a measure-matching question, an optimization problem on space of probability measures with the objective functional defined as a distance or divergence between probability measures, such as the Wasserstein distance, maximum mean discrepancy, and the Kullback-Leibler divergence. Note that $J$ and $N$ need not be equal, allowing for data upsampling or compression.
Example 3.9 (Generative Modeling).
Taking $d=2$, we consider the minimization of the following distance functional:
$G(\mu)=\int_{\mathbb{R}^{2}}\min_{x\in\mathbf{A}}|x-y|^{2}\,\mu(dy)-\frac{\alpha}{2}\int_{\mathbb{R}^{2}\times\mathbb{R}^{2}}\ln(|x-y|^{2})\,\mu(dx)\mu(dy),$
where $\alpha=3\times 10^{-9}$, the first term encourages the generated data to be close to the dataset $\mathbf{A}$, and the second term encourages the generated samples to spread out. In the numerical experiment reported in Figure 7, the target dataset $\mathbf{A}$ is a two-separate-horse silhouette consisting of 12,00012{,}000 points, while the generation is initialized from a snake-shaped configuration with 100,000100{,}000 points. We employ N=100, [^0] N=100{,}000 particles and discretize the time interval $[0,1]$ using M=10, [^0] M=10{,}000 time steps. The regularization parameter is set to $\varepsilon=10^{-15}$, and the conditional expectations in (3.11) are approximated with $S=100$ Monte Carlo samples. We first run a single outer iteration ($L=1$), and then recycle the resulting terminal particle configuration for two additional refinement iterations carried out over a shorter time horizon of length $0.001$ with time step $dt=10^{-6}$. The proposed SCM drives the particle system from the initial snake-shaped distribution toward an approximation of the target two-separate-horse distribution, achieving a high-fidelity reconstruction with an improved resolution; see Figure 7.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 7: Generative Modeling. The three frames of the first row showing the evolution of the particle distribution from initial shape of a snake (left) through two intermediate steps at $t=0.25$ (middle) and $t=0.95$ (right). The three frames of the second row showing the particle distribution at time $t=1.0$ (left), the refined particle distribution (middle) with two iterations, and the target empirical distribution (right) for comparison.
In view of the above example, the proposed SCM potentially provides a training-free and purely forward simulation-based alternative to diffusion models and Schrödinger bridge (SB) approaches for generative modeling. In contrast to diffusion models, which typically require an expensive offline training to learn a global score (see, e.g., [63]), SCM evaluates the optimal generative drift online through explicit probabilistic representations. Recent developments such as soft-constrained Schrödinger bridges [30, 46] alleviate the numerical instability associated with hard terminal constraints in classical SB by introducing penalty-based McKean-Vlasov formulations; however, these methods remain fundamentally rooted in potential-based constructions (e.g., via Doob’s $h$-transform) and often necessitate solving coupled forward-backward systems through global fixed-point iterations. By contrast, SCM avoids such forward-backward coupling. Indeed, by utilizing particle approximations together with the Cole-Hopf transformation, the Feynman-Kac representation, and the Bismut-Elworthy-Li formula, it yields an explicit representation of the generative drift and thereby reduces the computation to a single forward pass without training.
4 Conclusion
This paper develops a stochastic-control-based framework for global optimization by casting the original problem as the vanishing-regularization limit of a family of stochastic control problems. The objective functionals may be non-convex, non-differentiable, and defined on either finite-dimensional Euclidean spaces or the Wasserstein spaces of probability measures. This formulation enables the use of dynamic programming methods and, in finite dimensions, the Cole-Hopf transformation to linearize the associated HJB equation. The resulting representations via Feynman-Kac and the Bismut-Elworthy-Li formula yield derivative-free expressions for both value functions and optimal controls.
For optimization over $\mathbb{R}^{d}$, we establish convergence of the regularized control value to the global minimum at rate $\varepsilon\ln(1/\varepsilon)$ as $\varepsilon\to 0$. For optimization over Wasserstein space $\mathcal{P}_{2}(\mathbb{R}^{d})$, we introduce a regularized mean-field control formulation and approximate it by controlled $N$-particle systems. We prove quantitative convergence to the global optimum with error of order $\frac{C_{\varepsilon}}{N}+\varepsilon\ln\!\left(\frac{1}{\varepsilon}\right),$ as $N\to\infty$ and $\varepsilon\to 0$.
Based on the probabilistic representations, we further propose derivative-free numerical schemes and present computational experiments that support the theoretical findings. Reproducible code is available at: https://github.com/Jinnqiu/Stochastic-Control-Methods-for-Optimization. The numerical studies are intended as proof-of-concept illustrations rather than fully optimized implementations. A comprehensive computational study, including systematic parameter tuning, detailed benchmarking against established global optimization methods, and further applications in generative modeling and AI, is left for future work.